OpenStudy (torii897):

What is the solution to the equation 1 over 40 multiplied by x is equal to 1 over 8? A: x = 320 B:x = 5 C:x = 48 D:x = 32

OpenStudy (haseeb96):

x/40 =1/8 x= 5 is your correct answer

OpenStudy (torii897):

thank you (: